OpenStudy (anonymous):

a crazy apartment tenant throws a flower pot at a passing stranger 320 feet below on the street. assuming the flower pot is thrown with an initial velocity of 16 feet per second, how fast is it traveling when it (misses the stranger) an hits the street?

OpenStudy (anonymous):

and were supposed to use -32 for acceleration

OpenStudy (anonymous):

Vf^2=Vi^2+2a(Xf-Xi) Vf^2=(16)^2+2(9.8)(320) Vf=80.80ft/s downwards or -80.80ft/s

OpenStudy (anonymous):

oops. replace 32 for 9.8

OpenStudy (anonymous):

wow thanks a lot

OpenStudy (anonymous):

you get 144ft/s downwards

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

give me a medal. =D

OpenStudy (anonymous):

so why did you square the initial and final velocities? --ajain's friend

OpenStudy (anonymous):

that's one of the first formula's you're taught. you should be able to derive it from the position equation I believe, but I don't remember how to do that. =/

OpenStudy (anonymous):

oh thats ok i was just curious i figured it was just part of the equation